OpenStudy (anonymous):

I will give you a medal if can help me Solve the system using the elimination method. 5x+4y=-3 3x-7y=17

OpenStudy (anonymous):

please show work

OpenStudy (dumbcow):

Ok I would eliminate y because their coefficients have opposite signs LCM of 4 and 7 is 28 Multiply top equation by 7 Multiply bottom equation by 4 Then add equations 35x +28y = -21 12x -28y = 68 ----------------- 47x +0 = 47 47x = 47 x = 47/47 = 1 Substitute x=1 back into original equation to find y can you take it from here..